OpenStudy (anonymous):

How do you un-simplify a radical? Like: 3(squareroot)2

OpenStudy (anonymous):

3 squared is 9 bring the 9 inside the radial and multiply it by 2

OpenStudy (anonymous):

What about one that's like -5(squareroot)6

OpenStudy (anonymous):

same thing, 5 squared is 25, bring it inside the radial, multiply by 6 leave the minus sign out front

OpenStudy (anonymous):

THen how do you do a normal simplification like 3(square)40?

OpenStudy (anonymous):

*Square root

OpenStudy (anonymous):

\[\sqrt{40}=\sqrt{4\times 10}=\sqrt{4}\times \sqrt{10}=2\sqrt{10}\] look for perfect squares as factors of the number

OpenStudy (anonymous):

Thank you so much!